# SportsMetrics: Blending Text and Numerical Data to Understand Information Fusion in LLMs

### v1

@@ Abstract
[…]
@@ 1 Introduction
One of the most promising uses of LLMs is in handling a combination of unstructured texts and structured data. For example, determining if a patient can be discharged from the hospital may involve reviewing doctor notes, radiology and pathology reports, lab results, and other records that blend text and structured data Adams et al. (2021); Bardhan et al. (2022); Veen et al. (2023); Ben Abacha et al. (2023); LLM Assistants for online shopping need to process product catalogs, sales transactions, and customer queries Brynjolfsson et al. (2023); Loten (2023). However, summarizing key details from a mix of unstructured and structured sources remains a considerable challenge. An LLM must navigate text descriptions, link entities, aggregate numbers, handle discrepancies, and beyond.
@@
Information fusion focuses on synthesizing information from multiple textual sources to derive meaningful conclusions Barzilay et al. (1999). Current approaches involve summarizing multiple text documents, providing concise answers to user queries, and integrating summarization with natural language inference to deduce information Bhaskar et al. (2023); Caciularu et al. (2023); Sprague et al. (2022); Bostrom et al. (2022). The output is often a short text summary, the quality of which is difficult to evaluate Deutsch et al. (2021). Our approach differs by focusing on the numerical aspect of information fusion Geva et al. (2020). We enable the LLM to navigate through lengthy texts, gather crucial statistics, and develop a working memory to manage complex data queries.
@@
Our SportsMetrics benchmark presents three main benefits. First, it uses a broad range of sports data; they are dynamic narratives that LLMs cannot easily memorize. Second, it allows us to evaluate LLMs’ ability to track key statistics such as team points, assists, blocks, steals, and more, while also offering an overall game efficiency score for direct LLM comparison. Lastly, its use of widely understood sports terminology makes it more accessible to researchers than specialized medical language, making it an ideal benchmarking tool. While our current focus is on English, SportsMetrics also holds promise for multilingual applications.
@@ 2 Related Work
There is a growing need for a benchmark to evaluate LLMs’ information fusion capabilities, which offers clear, quantitative scores for comparing various LLMs. For example, Chatbot Arena Zheng et al. (2023) utilizes Elo ratings Boubdir et al. (2023), MT-Bench comprises of 80 multi-turn questions, and MMLU focuses on a model’s multitask accuracy across 57 tasks Hendrycks et al. (2021). Multi-document summarization offers a promising benchmark Huang et al. (2021); Wang et al. (2022); Xu et al. (2023). However, developing a summary scoring system poses challenges due to variables such as summary length, content coverage, and faithfulness Cao et al. (2022); Liu et al. (2023c); Krishna et al. (2023). Furthermore, we need to ensure that benchmark data has not been part of LLM pretraining datasets Liu et al. (2023c); Li et al. (2023). Sports game data, which combines static knowledge with player dynamics, presents an untapped opportunity for benchmarking LLMs.
@@
Our work relates to numerical reasoning, which uses arithmetic reasoning to tackle mathematical word problems. Prior datasets in this area include MathQA Amini et al. (2019), GSM8k Cobbe et al. (2021), SVAMP Patel et al. (2021), and MATH Liu et al. (2023d), targeting grade school level problems, and allowing models to generate answer explanations. The problems typically have brief descriptions, with the challenge lying in creating an expression tree and applying arithmetic knowledge. In contrast, our approach focuses on assessing LLMs’ ability to track key statistics across extremely long contexts.
@@ 3.3 Robustness Against Noise
Figure 5: Effective working memory is key in this task. While a sophisticated structure is possible, it also increases the likelihood of errors when populating values.

### v2

@@ Abstract
Yebowen Hu, Kaiqiang Song, Sangwoo Cho, Xiaoyang Wang, Hassan Foroosh, Dong Yu, Fei Liu§ University of Central Florida   Tencent AI Lab, Bellevue, WA   §Emory University {yebowen.hu, hassan.foroosh}@ucf.edu  {riversong, swcho, shawnxywang, dyu}@global.tencent.com fei.liu@emory.edu
@@ 1 Introduction
One of the most promising uses of LLMs is in handling a combination of unstructured texts and structured data. For example, determining if a patient can be discharged from the hospital may involve reviewing doctor notes, radiology and pathology reports, lab results, and other records that blend text and structured data Adams et al. (2021); Bardhan et al. (2022); Cai et al. (2022, 2023); Veen et al. (2023); Ben Abacha et al. (2023); LLM Assistants for online shopping need to process product catalogs, sales transactions, and customer queries Brynjolfsson et al. (2023); Loten (2023). Yet, summarizing key details from a mix of unstructured and structured sources remains a considerable challenge. An LLM must navigate text descriptions, link entities, aggregate numbers, handle discrepancies, and beyond.
@@
Information fusion focuses on synthesizing information from multiple textual sources to derive meaningful conclusions Barzilay et al. (1999). Current approaches involve summarizing multiple text documents, providing concise answers to user queries, and integrating summarization with natural language inference to deduce information Bhaskar et al. (2023); Caciularu et al. (2023); Sprague et al. (2022); Bostrom et al. (2022). The output is often a short text summary, the quality of which is difficult to evaluate Deutsch et al. (2021). In contrast, our approach emphasizes the numerical aspect of information fusion Geva et al. (2020); Zhu et al. (2021); Zhao et al. (2023); Reddy et al. (2024). We enable the LLM to navigate through lengthy texts, gather crucial statistics, and develop a working memory to manage complex data queries.
@@
Our SportsMetrics benchmark presents three main benefits. First, it leverages sports data, including team-player affiliations and play-by-play details; they are dynamic narratives that LLMs cannot easily memorize. Second, it allows us to evaluate LLMs’ ability to track key statistics such as team points, assists, blocks, steals, and more, while also offering an overall game efficiency score for direct LLM comparison. Lastly, its use of widely understood sports terminology makes it more accessible to researchers than specialized medical language, making it an ideal benchmarking tool. While our current focus is on English, SportsMetrics also holds promise for multilingual applications.
@@ 2 Related Work
There is a growing need for a benchmark to evaluate LLMs’ information fusion capabilities, which offers clear, quantitative scores for comparing various LLMs. For example, Chatbot Arena Zheng et al. (2023) utilizes Elo ratings Boubdir et al. (2023), MT-Bench comprises of 80 multi-turn questions, and MMLU focuses on a model’s multitask accuracy across 57 tasks Hendrycks et al. (2021). Multi-document summarization offers a promising benchmark Lebanoff et al. (2021); Huang et al. (2021); Wang et al. (2022); Xu et al. (2023). However, developing a summary scoring system poses challenges due to variables such as summary length, content coverage, and faithfulness Cao et al. (2022); Liu et al. (2023c); Krishna et al. (2023); Hu et al. (2023); Li et al. (2023); Xu et al. (2024); Joseph et al. (2024). Sports data, which combines static knowledge with player dynamics, presents an untapped opportunity for benchmarking LLMs.
@@
Our work relates to numerical reasoning, which uses arithmetic reasoning to tackle mathematical word problems. Prior datasets in this area include MathQA Amini et al. (2019), GSM8k Cobbe et al. (2021), SVAMP Patel et al. (2021), TAT-QA Zhu et al. (2021), FinQA Chen et al. (2022), MATH Liu et al. (2023d), DocMath-Eval Zhao et al. (2023), TABMWP Lu et al. (2023) and more, many allowing models to generate answer explanations. The problems typically have brief descriptions, with the challenge lying in creating an expression tree and applying arithmetic knowledge. In contrast, our approach focuses on assessing LLMs’ ability to track key statistics across extremely long contexts.
@@ 3.3 Robustness Against Noise
Figure 5: Effective working memory is key in this task. The variance in memory structure arises because we allowed each LLM to generate its JSON object as working memory, without enforcing a uniform schema. This step allows us to explore how each model organizes its memory to complete the task. Note that Claude’s ‘null’ values represent an initial state rather than an inability to aggregate information.
@@ 4 Experiments
Our results suggest that the difference in performance between GPT-3.5-Turbo-1106 and GPT-4 across basketball and football games stems from the scoring frequency in each sport. Basketball’s frequent scoring presents a challenge for GPT-4-1106-Preview to track all actions, while football, with less frequent scoring, is somewhat easier for the model to track. GPT-4-1106-Preview is optimized for handling extremely long contexts and it is less accurate in tracking frequent scoring. This distinct characteristic accounts for the varied performance of both models.
@@ 6 Limitations
Acknowledgements
We are grateful to the reviewers for their insightful feedback, which has helped enhance the quality of our paper. This research has been partially supported by the NSF CAREER award, #2303655.